OpenStudy (anonymous):

Provide an appropriate response.If a 13-inch wide rectangle is to approach the golden ratio, what should its length be? A. 34 in. B. 8 in. C. 20 in. D. 21 in.

OpenStudy (anonymous):

http://mathworld.wolfram.com/GoldenRatio.html 13*1.618 = 21.034 D.

OpenStudy (dumbcow):

it could also be 8 13/8 = 21/13 = 1.618

OpenStudy (dumbcow):

**approximately, and technically 21/13 is closer